OpenStudy (anonymous):

(HELP PLEASE) If a line passes through point (-1,2) and has slope m=a/b for a>0 and b>0, then which of the following points must lie of the graph of this line? A. (-1-a,2-b) B. (-1+b,2-a) C. (-1+b,2+a) D.(-1+a,2+b) I think it be between D or B. Im not sure really.

OpenStudy (loser66):

What is the equation of a line?

OpenStudy (anonymous):

y=mx+b?

OpenStudy (loser66):

yup, and you are given slope m = a/b, then y =?

OpenStudy (loser66):

oh, you confused , me too. Ok, this b is not the b in m = a/b,

OpenStudy (loser66):

I change a little bit: y = mx + h, ok?

OpenStudy (anonymous):

then y=a/bx+h?

OpenStudy (loser66):

m = a/b, hence \(y = \dfrac{a}{b}x +h\)

OpenStudy (loser66):

yup, now find h

OpenStudy (loser66):

do you know how?

OpenStudy (anonymous):

no...

OpenStudy (loser66):

you are given the point (-1, 2) that is x =-1, y =2, just plug it into y above.

OpenStudy (loser66):

\(h = y-\dfrac{a}{b}x\) replace the values of x and y to get h

OpenStudy (anonymous):

this is like extremely hard

OpenStudy (loser66):

so?? if you have another way, please help us. I know this way only :)

OpenStudy (anonymous):

I got it its C.

OpenStudy (loser66):

yup, me too

jimthompson5910 (jim_thompson5910):

The point (-1,2) is on this line the slope is m = a/b, so slope = rise/run = a/b rise = a run = b therefore you start at (-1,2) and go up 'a' units and go to the right 'b' units. You'll go from (-1,2) to (-1+b, 2+a)